A jetliner, traveling northward, is landing with a speed of 64.6 m/s. Once the jet touches down, it has 768 m of runway in which to reduce its speed to 14.6 m/s. Compute the average acceleration (magnitude and direction) of the plane during landing (take the direction of the plane's motion as positive).
Answer #1

. U = 64.6 m/s is s = 768 m v=14.6 m/s -utras (1406)2 = (64-6)2 + 2 (a ) (768 ) a = 3959.99 2(768) a = - 2. 578